OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m?

OpenStudy (anonymous):

x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2

OpenStudy (anonymous):